### v2

# HyperLoader: Integrating Hypernetwork-Based LoRA and Adapter Layers into Multi-Task Transformers for Sequence Labelling

Abstract
We present HyperLoader, a simple approach that combines different parameter-efficient fine-tuning methods in a multi-task setting. To achieve this goal, our model uses a hypernetwork to generate the weights of these modules based on the task, the transformer layer, and its position within this layer. Our method combines the benefits of multi-task learning by capturing the structure of all tasks while reducing the task interference problem by encapsulating the task-specific knowledge in the generated weights and the benefits of combining different parameter-efficient methods to outperform full-fine tuning. We provide empirical evidence that HyperLoader outperforms previous approaches in most datasets and obtains the best average performance across tasks in high-resource and low-resource scenarios.
HyperLoader: Integrating Hypernetwork-Based LoRA and Adapter Layers into Multi-Task Transformers for Sequence Labelling
Jesus-German Ortiz-Barajas1, Helena Gómez-Adorno2, Thamar Solorio1 1MBZUAI, 2IIMAS {jesus.ortizbarajas, thamar.solorio}@mbzuai.ac.ae helena.gomez@iimas.unam.mx
1 Introduction
Parameter-efficient fine-tuning techniques emerge as an alternative to conventional fine-tuning, where only a small number of parameters is updated to a downstream task Houlsby et al. (2019); Stickland and Murray (2019); Karimi Mahabadi et al. (2021a). These methods aim to achieve comparable performance to full fine-tuning by updating as few parameters as possible. However, a less studied research direction related to these methods is whether one can perform better than full fine-tuning with fewer parameters Mao et al. (2022).
Although these methods considerably reduce the number of parameters to achieve good performance on different tasks, a specialized model for each task remains necessary. Multi-task learning reduces the computational cost by training a single model while enabling information sharing through capturing the common structure underlying all tasks. Recently, researchers have developed approaches combining parameter-efficient fine-tuning techniques in multi-task settings Stickland and Murray (2019); Pfeiffer et al. (2020, 2021); Rücklé et al. (2021). Nevertheless, this approach can lead to underfitting in high-resource tasks and overfitting in low-resource tasks Lee et al. (2017). Another potential issue is task interference, where an improvement in the performance of one task reduces the performance of other tasks Wang et al. (2019).
An alternative to reduce the negative effects of multi-task learning is to use hypernetworks Ha et al. (2017) to generate separate weights for each task Karimi Mahabadi et al. (2021b); Ivison and Peters (2022). However, combining different parameter-efficient methods in a multi-task setting using hypernetworks is an unexplored research area. To address this limitation, we propose HyperLoader, a simple method that employs a neural network to generate the weights for a combination of two parameter-efficient fine-tuning techniques: adapters Houlsby et al. (2019) and LoRA Hu et al. (2021) based on the task, the transformer layer of the model and the position of the method within this layer. We use the encoder-decoder T5 model Raffel et al. (2020) for all experiments to take advantage of modelling the tasks as sequence-to-sequence tasks. We test our model in seven datasets from two Sequence Labelling tasks. The first task is Named Entity Recognition, a valuable tool in various real-world scenarios in the era of large language models such as healthcare and medical research Raza et al. (2022); Hu et al. (2024), Finance and Business Intelligence Zhou et al. (2023), and analyzing legal texts Trias et al. (2021). The second task is slot-filling, a crucial step to enable dialogue systems to accurately understand and fulfil user requests by extracting and organizing essential information from user inputs Cheng et al. (2023a, b); Firdaus et al. (2023).
Our model achieves the best average performance using the complete training and validation data for each task, as well as in a low-resource configuration with only 10% and 20% of the training and validation data available. We empirically demonstrate that the improvement is not only due to adding more trainable parameters but also by combining different parameter-efficient methods in a multi-task setting through hypernetworks, which adequately supports Sequence Labelling tasks in high- and low-resource scenarios.
Our main contributions are the following:
  • We propose a multi-task learning approach that combines different parameter-efficient fine-tuning methods based on hypernetworks conditioned on the task, the layer in the transformer model, and the position of the method within this layer.
  • We provide empirical results on different Sequence Labelling datasets demonstrating the effectiveness of our model compared to single-task and multi-task approaches using the entire datasets and in low-resource scenarios.
2 Related Work
Parameter-efficient fine-tuning techniques are an alternative to full fine-tuning by updating a small number of parameters per task, yielding a performance similar to that of full fine-tuning. One of the most popular approaches in this area is the adapter Houlsby et al. (2019), a small trainable bottleneck layer added in each transformer block of a model. It consists of down and up projections and has shown similar performance to full-fine-tuning. Other popular parameter-efficient fine-tuning approaches are Prefix-tuning Li and Liang (2021) and LoRA Hu et al. (2021). Prefix-tuning prepends a fixed-length, learnable sequence of prefix tokens to the input embeddings while the original model parameters remain unchanged. This approach optimizes a continuous prompt, effectively guiding the model’s behaviour for specific tasks with minimal computational overhead. LoRA adds trainable low-rank matrices and combines their outputs with the original matrices in the self-attention layer of the transformer; the main difference with the previous approaches is that it does not employ any activation function.
A less-studied area is how to achieve better performance than full-fine tuning using as few parameters as possible. The UniPELT framework Mao et al. (2022) addresses this limitation. This framework incorporates several parameter-efficient fine-tuning methods as sub-modules and learns how to activate them dynamically depending on the input or task setup using a gating mechanism, which assigns more weight to the sub-modules that increase the performance of a given task. They considered three different methods: Adapters, LoRA, and prefix-tuning, and they used the BERT base model for all their experiments.
Although promising, this approach still needs a specialized model for each task. The use of hypernetworks to generate the weights of parameter-efficient methods in a multi-task learning approach is a promising research area. By using this type of network, the model can capture the common structure underlying all target tasks and encapsulate the specific task knowledge by generating different weights based on the task, reducing the negative effects of this setting such as underfitting and overfitting in high-resource and low-resource tasks, respectively or task interference.
The HyperFormer++ model Karimi Mahabadi et al. (2021b) is the main foundation of our model. It uses task-conditioned hypernetworks to generate the weights of adapter layers and layer normalizations in a multi-task setting. The hypernetwork is trained to create adapter parameters specific to each task and layer, based on the embeddings of task and layer IDs. This hypernetwork is trained simultaneously for all tasks, allowing it to share knowledge across different tasks. At the same time, it reduces negative interference by producing distinct adapter layers for each individual task. The T5 model is used as the underlying model for all the experiments.
Another hypernetwork-based multi-task approach is Hyperdecoders Ivison and Peters (2022). This framework generates input-conditioned adapter layers for the decoder in an encoder-decoder model. This approach offers more flexibility by generating unique parameters for every input instead of using a learnt task embedding and allows making use of the similarities between samples across datasets.
Our HyperLoader model combines the benefits of both worlds. It uses different parameter-efficient fine-tuning techniques and task-conditional hypernetworks to generate its weights conditioned on the task, the transformer layer of the model, and the position of the parameter-efficient method within this layer. Our results show a performance improvement in all tasks and, on average, using the full training and validation data and simulating a low-resource setting.
3 Methodology
Multi-task hypernetwork-based approaches using parameter-efficient methods are a low-cost alternative for different NLP tasks that achieve comparable results to full-fine-tuning. Nevertheless, exploring approaches to obtain better results than updating all weights during fine-tuning is an unexplored research area. We introduce the HyperLoader model to address this gap. Our approach incorporates hypernetwork-based LoRA and Adapter layers into a multi-task transformer model, based on the HyperFormer++ model Karimi Mahabadi et al. (2021b).
Figure 1: Diagram of the HyperLoader model. The Adapter hypernetworks $h^{l}_{A_{D}}$ and $h^{l}_{A_{U}}$ produce the weights $D^{l}_{\tau}$ and $U^{l}_{\tau}$ for task-specific adapter modules. The LoRA hypernetworks $h^{l}_{LoRA_{A}}$ and $h^{l}_{LoRA_{B}}$ generate the $A$ and $B$ matrices for task-specific LoRA modules. Finally, the hypernetwork $h^{l}_{LN}$ creates the conditional layer normalization parameters $\beta_{\tau}$ and $\gamma_{\tau}$.
Our method enables the hypernetwork to generate the weights of the adapter layers and LoRA matrices based on the task, the transformer model’s layer, and the position of the adapter and LoRA matrix within this layer. Figure 1 shows the diagram of our proposed approach. We generate an input embedding $I_{\tau}$ based on the task $\tau$, the layer $l_{i}$ and the position of the Adapter and the LoRA matrices $p_{j}$. The embedding $I_{\tau}$ is the input to a set of hypernetworks ($h_{LN}^{l}$, $h^{l}_{A_{U}}$, $h^{l}_{A_{D}}$, $h^{l}_{LoRA_{B}}$, $h^{l}_{LoRA_{A}}$) to generate the weights for the adapter layers, the LoRA matrices and the layer normalization. Therefore, we only update the hypernetwork parameters, the input embedding projector $h^{{}^{\prime}}$ and the layer normalizations in $f_{\theta}(\cdot)$ during training, while the rest of parameters $\theta$ remain fixed.
In this section, we describe our approach in detail. First, we provide preliminary information by defining multi-task learning, which is the foundation of our work. Second, we describe the main components of our proposed HyperLoader model. Finally, we describe the format to convert Sequence Labelling into a sequence-to-sequence task.
3.1 Preliminaries: Multi-task learning
Our proposed model generates the weights of adapter layers and LoRA matrices based on the task, the transformer layer of the model, and the position of the adapter and the LoRA matrix in this layer in a multi-task setting. Given a set of tasks $\{D_{{\tau}_{\tau=1}^{T}\}}$ where $T$ is the total number of tasks, $D_{\tau}=\{(x_{\tau}^{i},y_{\tau}^{i})\}_{i=1}^{N_{\tau}}$ is the training data of the $\tau$-th task with $N_{\tau}$ instances and $f_{\theta}(\cdot)$ is a pre-trained model parameterized by $\theta$, the multi-task fine-tuning minimizes the loss shown in equation 1 on the training set, where $l$ is the cross-entropy loss and $w_{\tau}$ is the sampling weight for the $\tau$-th task.
$\mathcal{L}(\theta,\{D_{\tau}\}_{\tau=1}^{T})=\sum_{\tau=1}^{T}\sum_{(x_{\tau}^{i},y_{\tau}^{i})\in D_{\tau}}w_{\tau}l(f_{\theta}(x_{\tau}^{i}),y_{\tau}^{i})$ (1)
3.2 HyperLoader: Task Conditional Adapter Layers
Adapters Houlsby et al. (2019) are small sub-modules inserted within layers of a pre-trained transformer-based model before the skip connections. In the HyperLoader model, the conditional adapter $A^{l}$ for layer $l$ consists of a down-projection $\mathcal{D}^{l}_{\tau}\in\mathbb{R}^{h\times d}$ into a lower dimension $d_{bottleneck}$, a $GeLU$ non-linearity, and an up-projection $\mathcal{U}^{l}_{\tau}\in\mathbb{R}^{d\times h}$ that projects back into the original hidden layer dimension. The output of the adapter is defined in equation 2, where $x$ is the input hidden state, $LN^{l}_{\tau}$ is the task conditional layer normalization and the adapter weights $(\mathcal{D}^{l}_{\tau},\mathcal{U}^{l}_{\tau})$ are generated by a hypernetwork. We use a reduction factor $r=32$ for the down-projection in all the experiments.
$A^{l}_{\tau}(x)=LN^{l}_{\tau}\left(\mathcal{U}^{l}_{\tau}(GeLU(\mathcal{D}^{l}_{\tau}(x)))\right)+x$ (2)
3.3 HyperLoader: Task Conditional Layer Normalization
The task conditional layer normalization, defined in equation 3, involves element-wise multiplication denoted by $\odot$. Here, $\mu_{\tau}$ and $\sigma_{\tau}$ represent the mean and standard deviation of the training data for the $\tau$-th task, while $\gamma^{l}_{\tau}$ and $\beta^{l}_{\tau}$ are weights generated by a hypernetwork.
$LN^{l}_{\tau}(x^{i}_{\tau})=\gamma^{l}_{\tau}\odot\frac{x^{i}_{\tau}-\mu_{\tau}}{\sigma_{\tau}}+\beta^{l}_{\tau}$ (3)
3.4 HyperLoader: Task Conditional LoRA Matrices
Low-Rank Adaptation (LoRA) Hu et al. (2021) is a parameter-efficient fine-tuning technique that injects trainable rank decomposition matrices into each layer of the Transformer architecture. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, this method constrains its update by representing it with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$. The forward pass using LoRA is denoted in equation 4. Where $x$ is the input hidden state, $W_{0}$ is frozen and does not receive gradient updates and matrices $A$ and $B$ are generated by a hypernetwork. Based on previous work Hu et al. (2021); Mao et al. (2022), we only add Low-Rank matrices to the query and value matrices in the attention layers of the transformer. We use a low-dimensional rank $r=8$.
$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (4)
3.5 HyperLoader: Task Conditional Hypernetworks
A hypernetwork Ha et al. (2017) is a neural network that generates the parameters for another neural network. It captures shared information across tasks, while the generated task-specific adapters and layer normalization allow the model to adapt to each task individually, reducing negative interference in a multi-task setting. The input for these networks is a task embedding $z_{\tau}$ and its concatenation with a layer id embedding =$\mathcal{I}=\{l_{i}\}_{i=1}^{L}$ and an adapter position embedding $\mathcal{P}=\{p_{j}\}_{j=1}^{6}$. In the encoder-decoder model, we consider six positions for the parameter-efficient modules: the adapter layers and the LoRA matrices. We consider two options for the position of the adapter layers in each transformer block: after the self-attention layer or after the feed-forward layer. For LoRA matrices, we consider four positions: the A and B matrices in the self-attention layer and the A and B matrices in the cross-attention layer in the decoder of the model.
Using this configuration, the hypernetwork is able to generate different weights based on the task, position of the adapter, LoRA matrices and layer of the transformer network. The final input embedding $I_{\tau}$ for the hypernetwork is computed using a projector network $h^{{}^{\prime}}_{I}$ as shown in equation 5, where $z_{\tau}$, $l_{i}$ and $p_{j}$ are learnable parameters via back-propagation. We set the dimension of the task feature embedding $z_{\tau}$ to $d_{z_{\tau}}=512$. For the layer and position embeddings, we use a dimension $d_{l,p}=64$. Finally, for the final input embedding, we set a dimension of $d_{I_{\tau}}=64$.
$I_{\tau}=h^{{}^{\prime}}_{I}(z_{\tau},l_{i},p_{j})$ (5)
3.6 SentT’ format
Converting Sequence Labelling into a sequence-to-sequence task is essential to using an encoder-decoder model. We use the SentT’ Farina et al. (2023) format for the input and output sequences to achieve this goal. This format is obtained by turning a labelled sequence $S$ into a sequence of tokens $x_{1}\dots x_{L}$ of size $L$ using white-space splitting and then interleaving these tokens with the spacial sentinel tokens used to pre-train T5 to obtain the input string $S_{in}=s_{0}x_{1}s1\dots x_{L}s_{L}$. We use the simplified Beginning, Inside and Outside (sBIO) format to represent the output strings, where given a set $T$ of labels to annotate a text span, $t\in T$ is used to represent any token that starts a labelled span, a single tag $I$ for each token that continues a labelled span and $O$ to tag tokens that do not belong to labelled spans. Analogous to the input sequence, we interleave the tokenized output string with the sentinel tokens to obtain the output sequence using the SentT’ format $S_{out}=s_{0}t_{1}s_{1}\dots t_{L}s_{L}$. Table 1 shows an example from the SNIPS dataset of how data is converted into the SentT’ format.
Original text
play the song little robin redbreast
Encoded input using the SentT’ format
<extra_id_0> play <extra_id_1> the <extra_id_2> song <extra_id_3> little <extra_id_4> robin <extra_id_5> redbreast <extra_id_6>
Encoded output using the SentT’ format
<extra_id_0> O <extra_id_1> O <extra_id_2> MUSIC_ITEM <extra_id_3> TRACK <extra_id_4> I <extra_id_5> I <extra_id_6>
Table 1: Example of an input/output instance transformed into the SentT’ format. The text is tokenized using white-space splitting and then interleaving these tokens with the spacial sentinel tokens used to pre-train T5 to obtain the input string. Notice the use of the sBIO format to represent the output, where a label $t\in T$ is used to represent any token that starts a labelled span, a single tag $I$ for each token that continues a labelled span and $O$ to tag tokens that do not belong to labelled spans.
4 Experiments
This section describes our experiments for Sequence Labelling using the HyperLoader model. First, we introduce the datasets we used to train and evaluate our approach. Second, we provide the implementation, hyper-parameter and evaluation details of all experiments.
4.1 Corpora
We use seven publicly available corpora for Sequence Labelling tasks to cover different domains and distributions and gather robust empirical evidence. For slot-filling, we use four different dialogue-oriented datasets: ATIS Hemphill et al. (1990), SNIPS Coucke et al. (2018), and the English portions of two multilingual datasets, mTOP Li et al. (2021) and mTOP Schuster et al. (2019). For Named Entity Recognition, we use the MIT Corpora (Movie, MovieTrivia and Restaurant) [^1] Downloaded from https://groups.csail.mit.edu/sls/.
We used the original partitions for the mTOP and mTOD datasets, in the case of the ATIS and SNIPS datasets, we used the data partitions used originally by (Goo et al., 2018) [^2] Downloaded from https://github.com/MiuLab/SlotGated-SLU/tree/master/data, finally, we manually create a validation subset by sampling ten per cent of each original training partition for each MIT Corpus following previous methodologies Raman et al. (2022); Farina et al. (2023). Table 2 shows the statistics for the corpora we used in this work. It is important to mention that although there are some duplicates in the datasets, we do not remove these instances.
Corpora Train Val Test Labels
ATIS 4,478 500 893 83
SNIPS 13,084 700 700 39
MovieTrivia 7,034 782 1,953 12
Movie 8,797 978 2,443 12
Restaurant 6,894 766 1,521 8
mTOP 15,667 2,235 4,386 75
mTOD 30,521 4,181 8,621 9
Table 2: Statistics per partition of the used datasets. We use the training and validation partitions of Goo et al. (2018) for the ATIS and SNIPS dataset. For the MIT Corpora, we manually create the training and validation partitions following previous work Raman et al. (2022); Farina et al. (2023). Finally, we use the original training and validation partitions for the English portions of mTOP and mTOD.
4.2 Experimental details
We perform experiments to compare our model with other parameter-efficient fine-tuning approaches. The first goal is to observe the benefits of combining different parameter-efficient methods in a multi-task setting compared to a single-task setting. The second goal is to observe the benefits of the proposed model compared to other hypernetwork-based approaches in a multi-task setting.
We compare our approaches with the UniPELT framework Mao et al. (2022), HyperFormer model Karimi Mahabadi et al. (2021b) and Hyperdecoder model Ivison and Peters (2022) described in section 2. For all the implementations, we use the HuggingFace’s Transformers Wolf et al. (2020) and AdapterHub Poth et al. (2023) libraries.
We use the batch sizes, learning rates and epochs reported in the original paper of each baseline to run the experiments in our work. For our proposed HyperLoader model, we used the same codebase [^3] https://github.com/rabeehk/hyperformer (under the Apache License, Version 2.0) and almost the same experimental details described by Karimi Mahabadi et al. (2021b), that is, we employ the T5 base model, a batch size of 32, a constant learning rate of $3\times 10^{-4}$, $2^{18}$ steps in all experiments, save a checkpoint every 1,000 steps, and sample tasks with conventional temperature-based with temperature $T=10$ proportional to $p^{1/T}_{\tau}$, where $p_{\tau}=\frac{N_{\tau}}{\sum_{i=1}^{T}N_{\tau}}$ and $N_{\tau}$ is the number of training samples for the $\tau$-th task. The only difference in our work is that we choose the best model based on the loss value instead of the average evaluation metric performance. We performed the experiments on 4 Nvidia A100 SXM 40G GPUs.
4.3 Evaluation
We use the micro-averaged F1-score as the evaluation metric for all experiments following the CoNLL convention Tjong Kim Sang and De Meulder (2003), where an entity is considered correct only if the entity is predicted exactly as it appears in the gold data. We use the SeqEval framework Nakayama (2018) to compute the scores in all the experiments.
5 Results
This section presents the results of the experiments we performed in this work described in section 4.2.
5.1 Full dataset performance
Table 3 shows the performance of our proposed approach and the baselines using the 100% of training and evaluation data. The UniPELT framework follows a single-task fine-tuning approach, therefore, the trainable parameters correspond to each dataset considered. For this framework, we report the results using BERT as in the original paper and T5 to compare the performance depending on the model’s architecture. Initial experiments show that the use of the Prefix-tuning method and the gating mechanism decreases the performance of the encoder-decoder model; for this reason, we do not use these elements when applying UniPELT to T5.
Based on the results in Table 3, it is possible to observe that our proposed HyperLoader model outperforms the baselines in three of the seven datasets and obtains the highest average with 0.8811, followed by the HyperFormer model (0.8795), UniPELT with T5 (0.8740), the Hyperdecoder model (0.8709) and UniPELT with BERT (0.8537). In terms of single-task and multi-task settings, it is possible to observe that the combination of UNiPELT with T5 outperforms the other models in only one dataset (ATIS, with 0.9611); this result demonstrates that the use of hypernetworks improves the sharing information capacity of the models while reducing the negative transfer between tasks. Finally, when comparing the multi-task approaches based on hypernetworks, HyperLoader obtained a better average result than the HyperFormer model, which is the primary foundation of our work. In addition, since Hyperdecoder has more trainable parameters, the results indicate that the superior performance of our model is not only due to having more trainable parameters than the HyperFormer approach but also due to the combination of different parameter-efficient methods in a multi-task setting.
UniPELT UniPELT HyperFormer HyperLoader Hyperdecoder
Dataset BERT[110M/1.8M] $\clubsuit$ T5[220M/4.5M] $\clubsuit$ T5[228M/5M] T5[229M/6M] T5[239M/16M]
MovieTriva 0.6873 0.7210 0.7295 0.7140 0.7072
Movie 0.8500 0.8839 0.8788 0.8745 0.8700
Restaurant 0.7689 0.8039 0.8136 0.7965 0.7891
ATIS 0.9446 0.9611 0.9595 0.9599 0.9519
SNIPS 0.9238 0.9478 0.9462 0.9592 0.9377
mTOP 0.8432 0.8374 0.8638 0.8983 0.8752
mTOD 0.9580 0.9627 0.9648 0.9652 0.9649
Average 0.8537 0.8740 0.8795 0.8811 0.8709
Table 3: Performance of the HyperLoader model and the baselines using all train and evaluation examples. For each method, we indicate the used model and the total number of parameters / trainable parameters in brackets. $\clubsuit$ indicates a single-task fine-tuning approach; therefore, the number of trainable parameters is for each considered dataset. Bold fonts indicate the best result in each dataset and on average. Micro-averaged F1-score is reported.
5.2 Low-resource Setting Results
We evaluate the performance of our proposed model and the baselines in a low-resource setting to measure its robustness with a data scarcity constraint, a common scenario in NLP Hedderich et al. (2021). For this purpose, we down-sampled the training and validation partitions of all datasets to 10% and 20% of their original size and evaluated them in the full test subset. Table 4 shows the results of these experiments.
The first part of Table 4 contains the performance of the HyperLoader model and the baselines when only 10% of the training and validation data is used. It is evident that the combination of the UniPELT framework and the BERT model experiences the most significant performance decrease, from an average of 0.8537 to 0.6795. In contrast, the T5-based approaches show a smaller performance decrease in both single-task and multi-task settings. These results suggest that an encoder-decoder model is more capable of handling low-resource settings in both single-task and multi-task configurations. For the multi-task approaches, our HyperLoader model outperforms the HyperFormer and Hyperdecoder approaches in six out of seven datasets and on the average performance with a score of 0.8331, followed by HyperFormer (0.8262) and Hyperdecoder (0.8161).
The second part of Table 4 shows the results when 20% of the training and evaluation data is used. Using this amount of data our HyperLoader model outperforms the baselines in four of seven datasets and on the average performance. It is also noticeable that the combination of UniPELT and T5 using a single-task fine-tuning approach outperforms the other models only in the MovieTrivia dataset, with a micro-averaged F1-score of 0.7036 followed by HyperLoader with 0.6863, Hyperdecoder with 0.6767, HyperFormer with 0.6729 and finally UniPELT using BERT with 0.6105.
These results suggest that combining different parameter-efficient methods in a multi-task setting and generating their weights based on the task, the layer of the transformer model, and the position where the parameter-efficient method is applied, that is, the position of the Adapter layer or the LoRA matrix, is an effective method for Sequence Labelling tasks, yielding good performance using a parameter-efficient approach.
UniPELT UniPELT HyperFormer HyperLoader Hyperdecoder
Dataset BERT[110M/1.8M] $\clubsuit$ T5[220M/4.5M] $\clubsuit$ T5[228M/5M] T5[229M/6M] T5[239M/16M]
10% of Training and validation data
MovieTriva 0.5152 0.6717 0.6589 0.6695 0.6597
Movie 0.7809 0.8244 0.8311 0.8368 0.8202
Restaurant 0.5907 0.7049 0.7389 0.7499 0.7216
ATIS 0.5901 0.8919 0.9162 0.9172 0.9193
SNIPS 0.7398 0.8991 0.8943 0.8970 0.8727
mTOP 0.6048 0.7467 0.7976 0.8114 0.7760
mTOD 0.9351 0.9436 0.9467 0.9500 0.9436
Average 0.6795 0.8117 0.8262 0.8331 0.8162
20% of Training and validation data
MovieTriva 0.6105 0.7036 0.6729 0.6863 0.6767
Movie 0.8047 0.8444 0.8407 0.8546 0.8377
Restaurant 0.7114 0.7562 0.7517 0.7601 0.7401
ATIS 0.8048 0.9281 0.9382 0.9401 0.9311
SNIPS 0.8474 0.9235 0.9252 0.9238 0.9044
mTOP 0.7052 0.7863 0.8399 0.8474 0.8140
mTOD 0.9433 0.9522 0.9543 0.9541 0.9529
Average 0.7753 0.8420 0.8461 0.8523 0.8367
Table 4: Performance of the HyperLoader model and the baselines using 10% and 20% of the train and evaluation data. For each method, we indicate the used model and the total number of parameters / trainable parameters in brackets. $\clubsuit$ indicates a single-task fine-tuning approach; therefore, the number of trainable parameters is for each considered dataset. Bold fonts indicate the best result in each dataset and on average. Micro-averaged F1-score is reported.
Figure 2 illustrates the relationship between average performance and the percentage of trainable parameters per model. The symbol $\clubsuit$ denotes a single fine-tuning approach. It is important to note that, despite our model requiring more trainable parameters (6M) than the combination of UniPELT with T5 (4.5M), the percentage of parameters remains constant with increased training datasets.
Our proposed approach consistently outperforms other methods, using 10%, 20%, and 100% of the training and validation data. These results provide empirical evidence that combining different parameter-efficient fine-tuning methods in a multi-task setting is a robust strategy for low-resource scenarios. The performance improvement cannot be solely attributed to adding more parameters, as demonstrated by the hyperdecoder model having a more significant percentage of trainable parameters. Thus, the combination of parameter-efficient methods and weight generation emerges as a potent alternative for addressing Sequence Labelling tasks.
Figure 2: Average performance-percentage of trainable parameters plot using different portions of the datasets.$\clubsuit$ indicates a single-task fine-tuning approach
6 Conclusions
We proposed a method that combines parameter-efficient methods in a multi-task setting. Using a hypernetwork, we generate the weights of the adapter layers and LoRA matrices conditioned on the task, the transformer layer of the model, and the position of the adapter and LoRA matrices within this layer. Our method outperforms previous work in single-task and multi-task fine-tuning that combines different parameter-efficient methods and hypernetwork-based approaches using the full training and validation splits of each dataset and in a low-resource configuration using only 10% and 20% of those partitions to train the model. We provide empirical evidence that the improvement in performance is not only because of augmenting the trainable parameters since the hyperdecoder model has the largest number of trainable weights. Therefore, the combination of parameter-efficient methods and weight generation is a strong alternative to solving Sequence Labelling tasks in a multi-task setting.
7 Limitations and Future work
Although our model performs strongly compared to previous work in multi-task settings using a parameter-efficient fine-tuning approach that reduces negative transfer, underfitting, and overfitting across tasks, it still needs to have access to all datasets during training and requires complete re-training if a new task is added. Curriculum learning Bengio et al. (2009); Wang et al. (2021); Soviany et al. (2022); Piergiovanni et al. (2023) is a promising direction for future work and could alleviate the limitation above. This technique improves the model’s learning efficiency, often leading to faster convergence and better final performance.